OpenStudy (anonymous):

2. Perform this operation and report the answer to the correct number of significant figures: What is the mass of a cube of aluminum that is 4.0 cm on each edge? The density of aluminum is 2.7 g/cm3. Show all calculations leading to an answer.

OpenStudy (anonymous):

Perform this operation and report the answer to the correct number of significant figures: What is the mass of a cube of aluminum that is 4.0 cm on each edge? The density of aluminum is 2.7 g/cm3. Show all calculations leading to an answer. Vcube= length-width-height Vcube = 3.0cm.3.0cm.3.0cm=27cm3 2.7 g/cm3 = m/27cm3 m=2.7g/cm3.27cm3 m=73g

OpenStudy (cuanchi):

cube of aluminum that is 4.0 cm on each edge, no 3cm @ashleyrh1999 review the calculations